YoursTruly:

Solve by Completing the Square. 3x^2 + 20x + 12 = 0 x=20/3 and x=4 x=3 and x=-12 x=-2/3 and x=-6

YoursTruly:

@JustSaiyan @psirockin2

JustSaiyan:

I got this

JustSaiyan:

First, factor the left side: (3x+2)(x+6)=0 Set all factors equal to 0 3x+2=0 or x+6=0 \[x=-\frac{ 2 }{ 3 }~or~x=-6\] Answer: \[x=-\frac{ 2 }{ 3 }~or~x=-6\]

JustSaiyan:

So it is the final option. See? I can Math if I want.

YoursTruly:

"I can Math if I want."

JustSaiyan:

I can Algebra too.

YoursTruly:

k